OpenStudy (anonymous):

The difference of two positive numbers is 13. The larger is twice the smaller decreased by 7. Find the two numbers. (smaller value) (larger value)

OpenStudy (anonymous):

20 33

OpenStudy (anonymous):

it was 20 and 33. thanks though

OpenStudy (ybarrap):

ok so problem was l = 2s - 7 then.

OpenStudy (ybarrap):

d = 13 = l - s, where l = large number, s = small number l = 2s - 7 d = 2s - 7 - s=13 = s - 7 = 13 s = 13 + 7 = 20 l = 2*20 - 7 = 40 - 7 = 33 Ans, s = 20, l = 33

OpenStudy (anonymous):

thanks for showing how ya did it

OpenStudy (ybarrap):

np